OpenStudy (lgg23):

Calculus - Optimization Problem: A cylinder is inscribed in a right circular cone of height 5.5 and radius (at the base) equal to 3. What are the dimensions of such a cylinder which has maximum height? Radius = __?__ and Height = __?__

OpenStudy (anonymous):

Is there some info missing. A cylinder of max height would be 5.5 . there must be some other constraint like maximize the vol of the cylinder, etc

OpenStudy (lgg23):

I wish it had more info, but that is all that was given to me.

OpenStudy (mathmate):

which has maximum height? or maximum volume?

OpenStudy (mathmate):

If the question asks for maximum volume, then create the constraint for the cylinder. h/2 + r = 3 (h=0, r=3, h=6, r=0) The volume function is then V=pi r^2h = pi(r^2) 2(3-r) and dV/dr = 6pi r(2-r) = 0 for max or min. r=0 for min, and r=2 for maximum volume. Guess you can take care of the rest.